OpenStudy (anonymous):

find the unit tangent vector to the curve r(t)=((t^-1)+t),((t^-1)+t),-1) at t=1

OpenStudy (anonymous):

Find \(\mathbf{r}'(t)\) Can you do it?

OpenStudy (anonymous):

The Unit Tangent Vector is, obviously, a unit vector and it is defined as: \[=\frac{ r \prime (t) }{ \left| r \prime (t) \right| }\]

OpenStudy (anonymous):

\(\mathbf{r}'(t)\) is the tangent vector. To make it unit, divide by magnitude \[ \mathbf{\tau}(t) = \frac{\mathbf{r}'(t)}{\|\mathbf{r}'(t)\|} \]

OpenStudy (anonymous):

ok thanks i think i can do this now